OpenStudy (unknownunknown):

Session 6: Reading And Examples (Computing Determinants (PDF)). Example 1-1: What is this equation in reference to? 12 + (-8).... is this an error in the course material? I notice it again referenced in equation 4). How is this equation derived? Thanks.

OpenStudy (loser66):

Can you scan and post the PDF file?

OpenStudy (loser66):

I am with you. Let me read more but so far, you are right

OpenStudy (unknownunknown):

Interesting, thanks. I thought it was an error, but the same technique seems to be applied in Equation 4 as well.

OpenStudy (loser66):

yup, you are 100% right

OpenStudy (unknownunknown):

Do you think it's the same error in equation 4 too, at the bottom?

OpenStudy (loser66):

nope, the Laplace expansion is correct.

OpenStudy (loser66):

the first one, they use row 1 the second expansion, they use COLUMN 2

OpenStudy (loser66):

And for the sign of entry, it will be \[\left[\begin{matrix}+&-&+\\-&+&-\\+&-&+\end{matrix}\right]\]

OpenStudy (loser66):

hence for the first one, the numbers on Row 1 is +1, -0, +3 for the second one, when they use Column 2, the numbers and the signs will be -0, +2,-1 as shown

OpenStudy (loser66):

of course, you have to multiple the cofactor matrices on each of numbers. I explained why and how they use the method.

OpenStudy (unknownunknown):

I can see how the two end matrices are 2*(-7) -1*(-4) = 10, but the part under that where there is the equation.. |A| = −2 + 0 + 3 − 12 − 0 − (−1) = −10. How is that derived?

OpenStudy (unknownunknown):

-10*

OpenStudy (loser66):

ok, that is for second one, right?

OpenStudy (unknownunknown):

Yeah, I assume so.

OpenStudy (unknownunknown):

It says Checking By (1)?

OpenStudy (loser66):

\[2*\left|\begin{matrix}1&3\\2&-1\end{matrix}\right|=2*(1*(-1)-2*3)=2*(-7)\]

OpenStudy (unknownunknown):

Correct

OpenStudy (loser66):

so??

OpenStudy (unknownunknown):

Underneath that, the next sentence says something completely irrelevant? Checking by (1), and gives some weird equation also equalling -10, but the terms in it, I have no idea where they come from?

OpenStudy (loser66):

They might make mistake at the very first case on 2 x 2 matrix, but for the example 1-4, they are correct.

OpenStudy (unknownunknown):

But do you see that sentence there in the PDF, which begins with "Checking by (1)"?

OpenStudy (loser66):

Because they assume that it is correct.

OpenStudy (loser66):

nope,

OpenStudy (unknownunknown):

So that sentence is also incorrect there, the one appearing beneath equation 4 at the very bottom of the page?

OpenStudy (loser66):

(1) is correct, just only 1 line in example is not correct. Method (1) is CORRECT

OpenStudy (unknownunknown):

In the attachment I've highlighted it

OpenStudy (unknownunknown):

Is that circled red equation incorrect? I can't see how the -12 exists.

OpenStudy (loser66):

Can you post the previous PDF file? I want to see it. There might be something I didn't know. I would like to see the whole thing.

OpenStudy (loser66):

Give me more time, I will figure it out by checking from other site. I don't think they are wrong. MIT is an IVYleage school. They don't make that silly mistake, just our knowledge is not enough to understand.

OpenStudy (unknownunknown):

Yeah, I could understand a typo in the first equation, but another at the end of example 1.3? Seems a bit much. Thanks I really appreciate your help.

OpenStudy (unknownunknown):

It's late here so I'm going to go to bed, I'll accept your response But if you could post a reply with what you find out, I'd much appreciate it. Thanks again.

OpenStudy (loser66):

good night

OpenStudy (phi):

**** I can see how the two end matrices are 2*(-7) -1*(-4) = 10, but the part under that where there is the equation.. |A| = −2 + 0 + 3 − 12 − 0 − (−1) = −10. How is that derived? **** First, Example 1.1 is messed up. they present a 2x2 and show the work for a 3x3. For Example 1.3, they are using the expansion shown in Example 1.4 (multiplying "diagonals" and "antidiagonals" ) one mnemonic is to "extend" the matrix, and find the product of the diagonals, like this